OpenStudy (anonymous):

-4 times a number plus 16 is -116 find the number?

OpenStudy (welshfella):

-4x + 16 = -116 first add 16 to both sides then divide both sides by -4

OpenStudy (welshfella):

and you'll end up with x = something this will be your answer

OpenStudy (anonymous):

33?

OpenStudy (welshfella):

oh sorry silly me you SUBTRACT 16 from both sides

OpenStudy (welshfella):

-4x + 16 - 16 = -116-16 -4x = -132 x = 33 you were right

OpenStudy (anonymous):

its fine, i appreciate the help, its like 4:00 am and im just out of it

OpenStudy (welshfella):

lol ! I have no excuse, its 11 am in the UK...

OpenStudy (anonymous):

ah, im just gonna go to sleep nowww, thank you!!!

OpenStudy (anonymous):

well, go to sleep after I finish the rest of my hw

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

One thing though - the question is a bit ambiguous. The first part can be taken in another way It might be -4(x - 16) = -116 sorry gotta go

OpenStudy (anonymous):

yeh thx tho